OpenStudy (anonymous):

find the sum of the first 10 terms of the series 2-6+18-54+...

OpenStudy (asnaseer):

have you studied arithmetic and geometric progressions?

OpenStudy (asnaseer):

@kat762 if you have, then try and spot what type of series this is: 2, -6, 18, -54, ...

OpenStudy (anonymous):

you have an geometric series . The difference is a2 /a1 = -6/2 = -3 sn = a1 (1 - r^n)/(1 - r) n = 10 a1 = 2 (first term) r = -3